# Functionally Constrained Algorithm Solves Convex Simple Bilevel Problems

### v1

@@ 1 Introduction
$\min_{\mathbf{x}\in\mathbb{R}^{n}}\;f(\mathbf{x})\quad\text{s.t.}\quad\mathbf{x}\in{\mathcal{X}}_{g}^{*}\triangleq\mathop{\arg\min}_{\mathbf{z}\in{\mathbb{R}}^{n}}g(\mathbf{z}),$ (1)
where $f,g$ are convex and continuous functions. Such a problem is often referred to as “simple bilevel optimization” in the literature, as the upper-level objective function $f$ is minimized over the solution set of a lower-level problem. It captures a hierarchical structure and thus has many applications in machine learning, including lexicographic optimization [14, 12] and lifelong learning [17, 12]. Understanding the structure of simple bilevel optimization and designing efficient algorithms for it is vital and has gained massive attention in recent years [1, 5, 12, 26, 25, 11, 22, 13, 24, 23, 6, 10, 18].
@@
which we call a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution. Much progress has been achieved towards this goal  [10, 24, 23, 18, 22, 17]. We note all above algorithms fall in the class of zero-respecting algorithms (Assumption 3.4) and hence cannot obtain absolute optimal solutions.
@@
$\min_{{\bf{x}}\in\mathcal{B}({\bf{x}}_{0},D)}f({\bf{x}}),\quad{\rm s.t.}\quad\tilde{g}({\bf{x}})\triangleq g({\bf{x}})-\hat{g}^{*}\leq 0,$ (4)
where $\hat{g}^{*}$ is an approximation of $g^{*}$ such that $\hat{g}^{*}-g^{*}\leq\frac{\epsilon_{g}}{2}$, $D$ is an upper bound of initial distance $\|{\bf{x}}_{0}-{\bf{x}}^{*}\|_{2}$, and ${\mathcal{B}}({\bf{x}}_{0},D)$ is the Euclidean ball centered at the initial point ${\bf{x}}_{0}$ that contains the optimal solution ${\bf{x}}^{*}$ to Problem (1). Here we restrict the domain to $\mathcal{B}({\bf{x}}_{0},D)$ so that the subprocesses adopted in our method have a unified upper complexity bound (see Section 5.3). Then we further leverage the reformulation by Nesterov [19, Section 2.3.4] which relates the optimal value of Problem (4) to the minimal root of the following auxiliary function, where a discrete minimax problem defines the function value:
$\psi^{*}(t)=\min_{\mathbf{x}\in{\mathcal{B}}({\bf{x}}_{0},D)}\left\{\psi(t,{\bf{x}})\triangleq\max\left\{f({\bf{x}})-t,\tilde{g}({\bf{x}})\right\}\right\}.$ (5)
@@ 2 Related work
Kaushik and Yousefian [13] proposed the averaging iteratively regularized gradient (a-IRG) for convex optimization with variational inequality constraints, of which Problem (1) is a special case. a-IRG achieves the rate of ${\mathcal{O}}(1/\epsilon^{4})$. Shen et al. [24] proposed a method for solving Problem (1) with ${\mathcal{O}}(1/\epsilon^{3})$ complexity based on the online learning framework. When $f$ is Lipschitz and $g$ is smooth, Merchav and Sabach [18] proposed a gradient-based algorithm with ${\mathcal{O}}(1/\epsilon^{1/(1-\alpha)})$ complexity for any $\alpha\in(0.5,1)$. However, none of these methods can achieve the optimal rate of ${\mathcal{O}}(\epsilon^{-2})$.
@@
Samadi et al. [23] proposed the regularized accelerated proximal method (R-APM) with a complexity of ${\mathcal{O}}(1/\epsilon)$. Under the additional weak sharp minima condition on $g$, the complexity of R-APM improves to ${\mathcal{O}}\left(1/\sqrt{\epsilon}\right)$. However, this condition is often too strong and does not hold for many problems, Chen et al. [10] extended the result of [23] to the more general $\alpha$-Hölderian error bound condition. However, their method achieves the optimal rate only when $\alpha=1$, which reduces to the weak sharp minima condition. Jiang et al. [12] developed a Frank-Wolfe type algorithm (CG-BiO) with a complexity of $\mathcal{{\mathcal{O}}}\left(1/\epsilon\right)$, which approximates $\mathcal{X}_{g}^{*}$ by a cutting plane approach. Later on, Cao et al. [6] proposed an accelerated algorithm with a similar cutting plane approach to achieve the rate of ${\mathcal{O}}(1/\sqrt{\epsilon})$ under the weak sharp minima condition. Recently, Wang et al. [27] reduced Problem (1) to finding the smallest $c$ such that the optimal value of the following parametric problem is $g^{*}$: $\min_{{\bf{x}}\in{\mathbb{R}}^{n}}g({\bf{x}}),~{}{\rm s.t.}~{}f({\bf{x}})\leq c$. They adopted a bisection method to find such a $c$. To solve this parametric problem, Accelerated Proximal Gradient method is applied on $g$ with projection operator onto the sublevel set ${\mathcal{F}}_{c}=\{{\bf{x}}~{}|~{}f({\bf{x}})\leq c\}$, which we call sublevel set oracles. This leads to an upper bound of $\tilde{\mathcal{O}}(1/\sqrt{\epsilon})$. Such an oracle is obtainable for norm-like functions such as $f({\bf{x}})=\frac{1}{2}\|{\bf{x}}\|^{2}$. However, projection onto sublevel sets may be hard for more general functions, such as MSE loss or logistic loss. Compared with prior works [27, 6, 10, 23], our proposed methods achieve the $\tilde{\mathcal{O}}(1/\sqrt{\epsilon})$ rate under standard assumptions, without relying on the weak sharp minima condition or sublevel set oracles.
@@
Based on similar reformulation, Nesterov [19] have proposed algorithms for functionally constrained problems, of which Problem (4) is a special case: one for smooth problems in Section 2.3.5, and one for Lipschitz problems in Section 3.3.4. However, the first algorithm relies on the strong convexity of $f$ and $\tilde{g}$, which does not hold in our setups. Moreover, both algorithms require quadratic program (QP) sub-solvers, which are inefficient in large-scale problems: the first algorithm uses it to iteratively move $t$ to the root of $\psi^{*}(t)$, while the second algorithm uses it to project onto the sublevel set of a piecewise linear approximation of $\psi(t,{\bf{x}})$. In contrast, our proposed algorithms have the following advantages: (a) require no strong convexity assumption; (b) require no QP solvers that lead to additional computational cost; (c) have a unified framework for both smooth and Lipschitz problems.
@@ 3 Preliminaries
For any ${\bf{x}}\in{\mathbb{R}}^{n}$, let ${\bf{x}}_{[j]}$ represent the $j$-th coordinate of ${\bf{x}}$ for $j=1,\cdots,n$. The Euclidean ball centered at ${\bf{x}}$ with radius $D$ is denoted as $\mathcal{B}({\bf{x}},D)\triangleq\{{\bf{y}}\;|\;\|{\bf{y}}-{\bf{x}}\|_{2}\leq D\}$. For any closed convex set $\mathcal{C}\subseteq{\mathbb{R}}^{n}$, the Euclidean projection of ${\bf{x}}$ onto $\mathcal{C}$ is denoted by $\Pi_{\mathcal{C}}({\bf{x}})\triangleq\operatorname*{arg\,min}_{{\bf{y}}\in\mathcal{C}}\|{\bf{y}}-{\bf{x}}\|_{2}$. We say a function $h$ is $C$-Lipschitz in domain $\mathcal{Z}$ if $\|h({\bf{x}})-h({\bf{y}})\|_{2}\leq C\|{\bf{x}}-{\bf{y}}\|_{2}$ for all ${\bf{x}},{\bf{y}}\in\mathcal{Z}$. We say a differentiable real-valued function $h$ is $L$-smooth if it has $L$-Lipschitz continuous gradients.
We now state the assumptions on the function class required in our theoretical results. Complexity guarantees for finding an optimal solution to a convex optimization problem typically depend on the initial distance of $\|{\bf{x}}^{*}-{\bf{x}}_{0}\|_{2}$, where ${\bf{x}}_{0}$ is a pre-specified point as the initialization of algorithms. In our setting, we assume an explicit upper bound $D$ of this value.
@@
    For a pre-specified point ${\bf{x}}_{0}$, the initial distance $\|{\bf{x}}^{*}-{\bf{x}}_{0}\|_{2}$ is upper bounded by $D$, where ${\bf{x}}^{*}$ is the optimal solution to Problem (1).
@@
Consider Problem (1). We assume that $f$ and $g$ are $C_{f}$-Lipschitz and $C_{g}$-Lipschitz in $\mathcal{B}({\bf{x}}_{0},D)$ respectively. We call such problems $(C_{f},C_{g})$-Lipschitz problems.
@@
An iterative method $\mathcal{A}$ can access the objective functions $f$ and $g$ only through a first-order black-box oracle, which takes a test point $\hat{\bf{x}}$ as the input and returns $f(\mathbf{\hat{\bf{x}}}),g(\mathbf{\hat{\bf{x}}}),\partial f(\mathbf{\hat{\bf{x}}}),\partial g(\mathbf{\hat{\bf{x}}})$, where $\partial f(\hat{\bf{x}}),\partial g(\hat{\bf{x}})$ are arbitrary subgradients of the objective functions at $\hat{\bf{x}}$. $\mathcal{A}$ generates a sequence of test points $\{{\bf{x}}_{k}\}^{K}_{k=0}$ with
${\bf{x}}_{k}\in{\bf{x}}_{0}+\mathrm{Span}\{\partial f({\bf{x}}_{0}),\partial g({\bf{x}}_{0}),\cdots,\partial f({\bf{x}}_{k-1}),\partial g({\bf{x}}_{k-1})\}.$
This assumption generalizes the standard definition of zero-respecting algorithm for single-level minimization problems [19, 7]. Many existing first-order methods for Problem (1) fall into this algorithm class, including [10, 24, 23, 18, 22, 17].
@@
$\mathbf{x}_{k}\in\mathrm{Span}\{\nabla h(\mathbf{x}_{0}),\cdots,\nabla h(\mathbf{x}_{k-1})\},\;k\geq 1;\quad\mathbf{x}_{0}=\mathbf{0},$
@@
Definition 3.1 defines differentiable zero-chain functions. We can similarly define non-differentiable zero-chain, by requiring ${\bf{x}}_{k}$ to be in $\mathrm{Span}\{\partial h({\bf{x}}_{0}),\cdots,\partial h({\bf{x}}_{k-1})\}$ where $\partial h({\bf{x}})$ is a (possibly adversarial) subgradient of $h$ at ${\bf{x}}$.
@@ 5.1 Lower complexity bounds
We first establish the lower complexity bounds for finding a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution of $(L_{f},L_{g})$-smooth problems and $(C_{f},C_{g})$-Lipschitz problems.
@@
Given $L_{f},L_{g},D>0$. For any first-order algorithm $\mathcal{A}$ satisfying Assumption 3.4 and any initial point ${\bf{x}}_{0}$, there exists a $(L_{f},L_{g})$-smooth instance of Problem (1) such that the optimal solution ${\bf{x}}^{*}$ satisfies $\|{\bf{x}}^{*}-{\bf{x}}_{0}\|_{2}\leq D$ and $\mathcal{A}$ needs at least $\Omega\left(\max\left\{\sqrt{\frac{L_{f}}{\epsilon_{f}}},\sqrt{\frac{L_{g}}{\epsilon_{g}}}\right\}D\right)$ iterations to find a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution.
@@
Given $C_{f},C_{g},D>0$. For any first-order algorithm $\mathcal{A}$ satisfying Assumption 3.4 and any initial point ${\bf{x}}_{0}$, there exists a $(C_{f},C_{g})$-Lipschitz instance of Problem (1) such that the optimal solution ${\bf{x}}^{*}$ satisfies $\|{\bf{x}}^{*}-{\bf{x}}_{0}\|_{2}\leq D$ and $\mathcal{A}$ needs at least $\Omega\left(\max\left\{\frac{C_{f}^{2}}{\epsilon_{f}^{2}},\frac{C_{g}^{2}}{\epsilon_{g}^{2}}\right\}D^{2}\right)$ iterations to find a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution.
We prove these results by setting both $f({\bf{x}})$ and $g({\bf{x}})$ to be the worst-case zero-chain (Proposition A.1 and A.2), with $f({\bf{x}})$ depending only on the first half of the coordinates of ${\bf{x}}$ and $g({\bf{x}})$ depending only on the remaining coordinates. This construction decouples the optimization processes for $f$ and $g$. The full proof is provided in Appendix C.1.

### v2

@@ 1 Introduction
$\min_{\mathbf{x}\in\mathbb{R}^{n}}\;f(\mathbf{x})\quad\text{s.t.}\quad\mathbf{x}\in{\mathcal{X}}_{g}^{*}\triangleq\mathop{\arg\min}_{\mathbf{z}\in{\mathcal{Z}}}g(\mathbf{z}),$ (1)
where $f,g$ are convex and continuous functions and ${\mathcal{Z}}\subseteq\mathbb{R}^{n}$ is a compact convex set. Such a problem is often referred to as “simple bilevel optimization” in the literature, as the upper-level objective function $f$ is minimized over the solution set of a lower-level problem. It captures a hierarchical structure and thus has many applications in machine learning, including lexicographic optimization [14, 12] and lifelong learning [17, 12]. Understanding the structure of simple bilevel optimization and designing efficient algorithms for it is vital and has gained massive attention in recent years [1, 5, 12, 26, 25, 11, 22, 13, 24, 23, 6, 10, 18].
@@
which we call a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution. Much progress has been achieved towards this goal  [10, 24, 23, 18, 22, 17]. We note all the above algorithms fall in the class of zero-respecting algorithms (Assumption 3.4) and hence cannot obtain absolute optimal solutions.
@@
$\min_{{\bf{x}}\in{\mathcal{Z}}}f({\bf{x}}),\quad{\rm s.t.}\quad\tilde{g}({\bf{x}})\triangleq g({\bf{x}})-\hat{g}^{*}\leq 0,$ (4)
Then we further leverage the reformulation by Nesterov [19, Section 2.3.4] which relates the optimal value of Problem (4) to the minimal root of the following auxiliary function, where a discrete minimax problem defines the function value:
$\psi^{*}(t)=\min_{{\mathcal{Z}}}\left\{\psi(t,{\bf{x}})\triangleq\max\left\{f({\bf{x}})-t,\tilde{g}({\bf{x}})\right\}\right\}.$ (5)
@@ 2 Related work
Kaushik and Yousefian [13] proposed the averaging iteratively regularized gradient method (a-IRG) for convex optimization with variational inequality constraints, of which Problem (1) is a special case. a-IRG achieves the rate of ${\mathcal{O}}(1/\epsilon^{4})$. Shen et al. [24] proposed a method for solving Problem (1) with ${\mathcal{O}}(1/\epsilon^{3})$ complexity based on the online learning framework. When $f$ is Lipschitz and $g$ is smooth, Merchav and Sabach [18] proposed a gradient-based algorithm with ${\mathcal{O}}(1/\epsilon^{1/(1-\alpha)})$ complexity for any $\alpha\in(0.5,1)$. However, none of these methods can achieve the optimal rate of ${\mathcal{O}}(\epsilon^{-2})$.
@@
Samadi et al. [23] proposed the regularized accelerated proximal method (R-APM) with a complexity of ${\mathcal{O}}(1/\epsilon)$. Under the additional weak sharp minima condition on $g$, the complexity of R-APM improves to ${\mathcal{O}}\left(1/\sqrt{\epsilon}\right)$. However, this condition is often too strong and does not hold for many problems. Chen et al. [10] extended the result of [23] to the more general $\alpha$-Hölderian error bound condition. However, their method achieves the optimal rate only when $\alpha=1$, which reduces to the weak sharp minima condition. Jiang et al. [12] developed a conditional gradient type algorithm (CG-BiO) with a complexity of $\mathcal{{\mathcal{O}}}\left(1/\epsilon\right)$, which approximates $\mathcal{X}_{g}^{*}$ similar to the cutting plane approach. Later on, Cao et al. [6] proposed an accelerated algorithm with a similar cutting plane approach to achieve the rate of ${\mathcal{O}}(1/\sqrt{\epsilon})$ under the weak sharp minima condition. Recently, Wang et al. [27] reduced Problem (1) to finding the smallest $c$ such that the optimal value of the following parametric problem is $g^{*}$: $\min_{{\bf{x}}\in{\mathbb{R}}^{n}}g({\bf{x}}),~{}{\rm s.t.}~{}f({\bf{x}})\leq c$. They adopted a bisection method to find such a $c$. To solve this parametric problem, Accelerated Proximal Gradient method is applied on $g$ with projection operator onto the sublevel set ${\mathcal{F}}_{c}=\{{\bf{x}}~{}|~{}f({\bf{x}})\leq c\}$, which we call sublevel set oracles, which we call sublevel set oracles. This leads to an upper bound of $\tilde{\mathcal{O}}(1/\sqrt{\epsilon})$. Such an oracle is obtainable for norm-like functions such as $f({\bf{x}})=\frac{1}{2}\|{\bf{x}}\|^{2}$. However, projection onto sublevel sets may be hard for more general functions, such as MSE loss or logistic loss. Compared with prior works [27, 6, 10, 23], our proposed methods achieve the $\tilde{\mathcal{O}}(1/\sqrt{\epsilon})$ rate under standard assumptions, without relying on the weak sharp minima condition or the sublevel set oracles.
@@
Based on similar reformulation, Nesterov [19] has proposed algorithms for functionally constrained problems, of which Problem (4) is a special case: one for smooth problems in Section 2.3.5, and one for Lipschitz problems in Section 3.3.4. However, the first algorithm relies on the strong convexity of $f$ and $\tilde{g}$, which does not hold in our setups. Moreover, both algorithms require quadratic program (QP) sub-solvers, which are inefficient in large-scale problems: the first algorithm uses it to iteratively move $t$ to the root of $\psi^{*}(t)$, while the second algorithm uses it to project onto the sublevel set of a piecewise linear approximation of $\psi(t,{\bf{x}})$. In contrast, our proposed algorithms have the following advantages: (a) require no strong convexity assumption; (b) require no QP solvers that lead to additional computational cost; (c) have a unified framework for both smooth and Lipschitz problems.
@@ 3 Preliminaries
For any ${\bf{x}}\in{\mathbb{R}}^{n}$, let ${\bf{x}}_{[j]}$ represent the $j$-th coordinate of ${\bf{x}}$ for $j=1,\cdots,n$. We use $\text{supp}({\bf{x}}):=\{j\in[d]~{}:~{}{\bf{x}}_{[j]}\not=0\}$ to denote the support of ${\bf{x}}$. The Euclidean ball centered at ${\bf{x}}$ with radius $R$ is denoted as $\mathcal{B}({\bf{x}},R)\triangleq\{{\bf{y}}\;|\;\|{\bf{y}}-{\bf{x}}\|_{2}\leq R\}$. For any closed convex set $\mathcal{C}\subseteq{\mathbb{R}}^{n}$, the Euclidean projection of ${\bf{x}}$ onto $\mathcal{C}$ is denoted by $\Pi_{\mathcal{C}}({\bf{x}})\triangleq\operatorname*{arg\,min}_{{\bf{y}}\in\mathcal{C}}\|{\bf{y}}-{\bf{x}}\|_{2}$. We say a function $h$ is $C$-Lipschitz in domain $\mathcal{Z}$ if $\|h({\bf{x}})-h({\bf{y}})\|_{2}\leq C\|{\bf{x}}-{\bf{y}}\|_{2}$ for all ${\bf{x}},{\bf{y}}\in\mathcal{Z}$. We say a differentiable real-valued function $h$ is $L$-smooth if it has $L$-Lipschitz continuous gradients.
We now state the assumptions required in our theoretical results.
@@
    The feasible set ${\mathcal{Z}}$ is convex and compact with radius $D=\sup_{{\bf{x}},{\bf{y}}\in{\mathcal{Z}}}\|{\bf{x}}-{\bf{y}}\|_{2}$.
The compactness assumption ensures that the subprocesses adopted in our method have a unified upper complexity bound (see Section 5.3). We note that other works involving bisection procedures, such as Wang et al. [27], may also need to address this issue to derive an explicit dependence on the distance term (although they did not state it formally in their paper). For unconstrained problems, if we know that the initial distance $\|{\bf{x}}^{*}-{\bf{x}}_{0}\|_{2}$ is upper bounded by $R$, we can simply take ${\mathcal{Z}}={\mathcal{B}}({\bf{x}}_{0},R)$.
@@
Consider Problem (1). We assume that $f$ and $g$ are $C_{f}$-Lipschitz and $C_{g}$-Lipschitz in ${\mathcal{Z}}$ respectively. We call such problems $(C_{f},C_{g})$-Lipschitz problems.
@@
An iterative method $\mathcal{A}$ can access the objective functions $f$ and $g$ only through a first-order black-box oracle, which takes a test point $\hat{\bf{x}}$ as the input and returns $\partial f(\mathbf{\hat{\bf{x}}}),\partial g(\mathbf{\hat{\bf{x}}})$, where $\partial f(\hat{\bf{x}}),\partial g(\hat{\bf{x}})$ are arbitrary subgradients of the objective functions at $\hat{\bf{x}}$. $\mathcal{A}$ generates a sequence of test points $\{{\bf{x}}_{k}\}^{K}_{k=0}$ with
${\rm supp}({\bf{x}}_{t+1})\subseteq{\rm supp}({\bf{x}}_{0})\cup\left(\bigcup_{0\leq s\leq t}{\rm supp}\left(\partial f({\bf{x}}_{s})\right)\cup{\rm supp}(\partial g({\bf{x}}_{s}))\right).$ (6)
This assumption generalizes the standard definition of zero-respecting algorithm for single-level minimization problems [19, 7]. Many existing methods that incorporate a gradient step in the update for Problem (1) clearly fall within this class of algorithms, including those proposed by [10, 24, 23, 18, 22, 17], since the gradient step ensures that ${\bf{x}}_{k}\in{\bf{x}}_{0}+\mathrm{Span}\{\partial f({\bf{x}}_{0}),\partial g({\bf{x}}_{0}),\dots,\partial f({\bf{x}}_{k-1}),\partial g({\bf{x}}_{k-1})\}$. In the appendix, we show that the proposed algorithm in this paper and the conditional gradient type methods [12, 5] also satisfy the condition (6) when the domain is a Euclidean ball centered at ${\bf{x}}_{0}$ (see Proposition C.1 and Remark C.1).
@@
${\rm supp}({\bf{x}}_{t+1})\subseteq\bigcup_{0\leq s\leq t}{\rm supp}\left(\nabla h({\bf{x}}_{s})\right),\;k\geq 1;\quad\mathbf{x}_{0}=\mathbf{0},$
@@
Definition 3.1 defines differentiable zero-chain functions. We can similarly define non-differentiable zero-chain, by requiring ${\rm supp}({\bf{x}}_{t+1})$ to be in $\bigcup_{0\leq s\leq t}{\rm supp}\left(\partial h({\bf{x}}_{s})\right)$, where $\partial h({\bf{x}})$ is a (possibly adversarial) subgradient of $h$ at ${\bf{x}}$.
@@ 4 Finding absolute optimal solutions is hard
The constructions of the above hardness results are motivated by the work [9], which demonstrated that for general bilevel optimization problems of the form $\min_{{\bf{x}}\in{\mathbb{R}}^{n},{\bf{y}}\in{\mathbb{R}}^{m}}f({\bf{x}},{\bf{y}})$ subject to ${\bf{y}}\in\arg\min_{{\bf{z}}\in{\mathbb{R}}^{m}}g({\bf{x}},{\bf{z}})$, there exists a "hard instance" in which any zero-respecting algorithm always yields ${\bf{x}}_{k}={\bf{x}}_{0}$ for all $1\leq k\leq T$. Although our construction has a very similar high-level idea to [9], the $f$ and $g$ we construct are different from the functions in [9] since our desired conclusion is different.
@@ 5.1 Lower complexity bounds
We first establish the lower complexity bounds for finding a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution of $(L_{f},L_{g})$-smooth problems and $(C_{f},C_{g})$-Lipschitz problems. The results follow directly from existing lower bounds for single-level optimization problems, as simple bilevel optimization is a more general framework. Although the proof is straightforward, we present the results because establishing a precise lower bound is essential for demonstrating that an algorithm is truly near-optimal.
@@
Given $L_{f},L_{g},D>0$. For any first-order algorithm $\mathcal{A}$ satisfying Assumption 3.4 and any initial point ${\bf{x}}_{0}$, there exists a $(L_{f},L_{g})$-smooth instance of Problem (1) on the domain ${\mathcal{Z}}=\mathcal{B}({\bf{x}}_{0},D)$ such that the optimal solution ${\bf{x}}^{*}$ is contained in ${\mathcal{Z}}$ and $\mathcal{A}$ needs at least $\Omega\left(\max\left\{\sqrt{\frac{L_{f}}{\epsilon_{f}}},\sqrt{\frac{L_{g}}{\epsilon_{g}}}\right\}D\right)$ iterations to find a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution.
@@
Given $C_{f},C_{g},D>0$. For any first-order algorithm $\mathcal{A}$ satisfying Assumption 3.4 and any initial point ${\bf{x}}_{0}$, there exists a $(C_{f},C_{g})$-Lipschitz instance of Problem (1) on the domain ${\mathcal{Z}}=\mathcal{B}({\bf{x}}_{0},D)$ such that the optimal solution ${\bf{x}}^{*}$ is contained in ${\mathcal{Z}}$ and $\mathcal{A}$ needs at least $\Omega\left(\max\left\{\frac{C_{f}^{2}}{\epsilon_{f}^{2}},\frac{C_{g}^{2}}{\epsilon_{g}^{2}}\right\}D^{2}\right)$ iterations to find a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution.